# Gender Dynamics and Homophily in a Social Network of LLM Agents

### v1

@@ Appendix A Data collection and preprocessing
Table S3: Results of gender-score calculation for a subset of a Twitter dataset.
From Table S3, we observe that the model achieved an overall accuracy of 0.84 in both runs. For consistency, we found that 97.4% of profiles (974 out of 1000) were assigned the same gender category in both runs, and 85.4% (854 out of 1000) received the same gender score. These results highlight the robustness of the model in consistently identifying gender across multiple runs.
@@
We use the followership data to construct a sequence of directed networks that summarise the the social structure of the Chirper community over the year. Nodes correspond to Chirpers and a directed edge from $i$ to $j$ indicates that $i$ follows $j$.
@@
Each node in week $t$ is annotated with the gender-score $X_{it}$ inferred from its content in that week, when available. If no score is available, meaning that the agent appears in the followership network in week $t$ but has not posted any content in that week, so that $X_{it}$ is missing, we impute its value using the closest available week in time. To do this, we take the most recent previous week when possible and otherwise the nearest subsequent week. These weekly networks, together with the associated gender-scores, form the basis for our analysis of gender dynamics, homophily, and the separation between selection and influence in later sections.
Appendix B Null Networks
@@
For every directed weekly follow graph, we construct $100$ realisations using a directed configuration model that preserves each node’s in-degree and out-degree sequence. Concretely, for week $t$, we draw a directed multigraph from the observed in and out degree sequences, convert it to a simple directed graph by collapsing parallel edges, remove self-loops, relabel nodes back to their original identities, and retain their weekly gender scores $X_{it}$. This yields null networks that match empirical degree heterogeneity while randomising ties otherwise.
@@
We generate an additional $100$ realisations per week using a directed joint degree model, which preserves not only the in and out degree sequences but also the joint distribution of source out-degrees and target in-degrees across edges (the $n_{kk^{\prime}}$ mixing matrix), that is, degree–degree mixing. After removing self-loops and duplicate edges from the empirical weekly graph to estimate $n_{kk^{\prime}}$, we sample null graphs from this joint degree ensemble. We then assign gender scores to null nodes by drawing from the empirical distribution of $X_{it}$ within each $(k_{\mathrm{out}},k_{\mathrm{in}})$ degree class, thereby maintaining any attribute?degree relationship present in the data without imposing additional similarity.
@@
Appendix C STERGM implementation details
@@
Edges are marked as active for the duration of their first appearance in the block to the end of the block. In particular, if a directed tie $i\to j$ is first observed in week $t_{0}$ within the block, it is treated as active on $t_{0},T+1)$, where $T$ is the last week in the block. Because follow ties are persistent in our data (we do not observe unfollow events), we do not model dissolution, and the STERGM is fitted in formation-only mode.
@@
The construction of the global roster and weekly gender score panel mirrors the ERGM analysis (Appendix LABEL:app:ergm-details). Starting from all 52 weekly edge lists, we recover the weekly gender scores $X_{it}$ for each actor week pair, fill short temporal gaps via last observation carried forward and next observation carried backward, and retain only actors who have at least one observed value. This yields a cleaned global roster and a completed weekly panel of gender scores.
Within each block, we restrict attention to actors who appear in at least one edge within the block and intersect this block-specific roster with the cleaned global roster. For each week $t$ in the block, we then extract the completed scores for these actors and standardise them. The resulting standardised scores are attached as a time varying vertex attribute score on the interval $[t,t+1)$ for all actors present in that week.
@@
On each dynamic network, we fit a formation-only STERGM using the tergm package in R [28, 29]. Let $A^{(t)}$ and $A^{(t+1)}$ denote the adjacency matrices at the start and end of an interval, and let $X_{t}$ be the vector of standardised gender-scores at time $t$. The formation model specifies, for dyads with $A^{(t)}_{ij}=0$, the conditional probability that a new tie $i\to j$ is formed by $t+1$ as an exponential-family model whose sufficient statistics include an edge formation count, a mutual formation count, and a measure of gender dissimilarity on newly formed ties.
In compact form, the formation component uses the same three structural effects as the cross-sectional ERGMs,
@@
where $s_{\mathrm{edges}}$ controls for baseline formation density, $s_{\mathrm{mutual}}$ captures the tendency to form reciprocal ties, and $s_{\mathrm{abs}}$ accumulates the absolute difference in standardised gender-scores across the dyads on which formation is being considered. The corresponding formation parameter vector $\phi=(\phi_{\mathrm{edges}},\phi_{\mathrm{mutual}},\phi_{\mathrm{abs}})^{\top}$ governs the log odds of a new tie forming between weeks $t$ and $t+1$. The coefficient $\phi_{\mathrm{abs}}$ is the primary quantity of interest, with negative values indicating gender based homophily in the formation of new follow ties and positive values indicating heterophily. As in the ERGM analysis, we report $\exp(\hat{\phi}_{\mathrm{abs}})$ as the odds ratio for a one-standard-deviation increase in the absolute difference in gendered performance between two actors.
For computational diagnostics, we first fit formation-only STERGMs using conditional maximum likelihood estimation with inner models restricted to the maximum pseudo-likelihood (MPLE) step. We then use the default conditional maximum likelihood estimation in tergm, which refines the initial estimates via Markov chain Monte Carlo.
Appendix D Panel regression implementation details
@@
We work with a weekly panel covering weeks $t=1,\dots,52$. For each week, we load the annotated edge list, where each row records a directed follow event from a sender to a receiver, along with their weekly gender scores. As in the ERGM and STERGM analyses, we first build a long table of triples $(\text{week},\text{id},\text{gender-score})$ by pooling the sender and receiver columns and using the unique weekly gender score for each agent.
To obtain a complete weekly series of gender scores for each agent, we apply the same forward- and backward-time imputations as in the ERGM construction. Within each agent, we carry the last observed value forward to fill short gaps, and then use the next observed value to fill any remaining gaps. Actors who never have an observed gender score in any week are removed from the panel.
Using the followership data, we then compute for each agent $i$ and week $t$ the average gender-score of their followees in that week, denoted $\bar{X}_{N(i),t}$. This is obtained by averaging the weekly gender-scores of all agents followed by $i$ in week $t$.
For each agent $i$ and week $t\geq 2$ we construct the lagged gender-score $X_{i,t-1}$, the lagged mean followee score $\bar{X}_{N(i),t-1}$, and the weekly change
@@
In the final estimation sample we standardise $\Delta X_{it}$, $X_{i,t-1}$ and $\bar{X}_{N(i),t-1}$ to have mean zero and unit variance across the sample, so that coefficients can be interpreted in units of standard deviations.
@@
Our main influence model relates weekly changes in gender performance to the previous week’s own score and the previous week’s mean score of followees:
@@
We first estimate this specification using ordinary least squares (OLS) with agent and week fixed effects, clustering standard errors at the agent level. In this fixed effects OLS model, $\varphi$ measures self-correction, that is, how much an agent’s change in gender-score depends on their own previous level after controlling for time-invariant agent heterogeneity and common weekly shocks. The coefficient $\gamma_{\mathrm{inf}}$ measures peer alignment, describing how much an agent’s change depends on the previous week’s mean gender score of their followees. Because all variables are standardised, $\varphi$ and $\gamma_{\mathrm{inf}}$ can be read as changes in $\Delta X_{it}$ in standard deviation units associated with a one standard deviation change in the corresponding predictor.
@@
As an instrument for $\bar{X}_{N(i),t-1}$, we use the mean lagged gender-score of followees of followees, that is, agents at distance two in the follow network at time t1t-1. For each week $t$ and ego $i$ we consider all paths of length two of the form
@@
present in the week t1t-1 follow network, where $j$ is a followee of $i$ and $k$ is a followee of $j$. We exclude cases where $k=i$ or where $k$ is already a direct followee of $i$, so the resulting set consists of friends of friends who are not direct neighbours. We then take the mean of their lagged gender-scores and use this as an instrument $Z_{it}$ for $\bar{X}_{N(i),t-1}$.
In the first stage we regress the lagged mean followee score on the instrument, including the same agent and week fixed effects as in the outcome equation.
@@
so that $\gamma_{\mathrm{inf}}$ is an instrumental variables estimate of the causal effect of peers on changes in manifested gender, under the assumption that friends of friends affect the ego’s change only through their influence on the behaviour of direct followees.
We use the fixed-effects OLS estimates to describe overall correlations between peer and ego behaviour and the fixed-effects IV estimates to quantify causal peer influence. A one standard deviation increase in the lagged mean gender-score of followees is interpreted as producing an estimated $\gamma_{\mathrm{inf}}$ standard deviation change in the ego’s weekly gender-score, holding own previous score, agent effects, and week effects constant.
Full year and windowed estimators
We first estimate both the fixed effects OLS and fixed effects IV models on the full year panel using all weeks $t=2,\dots,52$ for which lagged variables are defined. To examine how the strength of influence evolves over time, we then repeat the same estimation on consecutive eight-week windows, always including an additional look-back week to define lags. For each window, we record the estimates of $\varphi$ and $\gamma_{\mathrm{inf}}$ from both OLS and IV, together with standard weak instrument diagnostics such as the first stage $F$ statistic and the Wu?Hausman test for endogeneity of the peer term.
Appendix E Structural network measures: detailed definitions
This appendix provides the formal definitions and computational steps for the structural measures reported in Section (i). All measures are computed separately for each weekly followership network. Let $A^{(t)}$ denote the directed adjacency matrix for week $t$, and let $n_{t}$ be the number of active agents in that week.
@@ Appendix E Structural network measures: detailed definitions
Social networks are typically sparse, meaning that only a small fraction of all possible ties are present. Tracking density over time reveals how quickly the network becomes more interconnected as agents form follower relationships. Network density measures the proportion of realised ties out of all possible directed ties and is defined as:
@@
Appendix F Additional structural network plots
Figure S1: Evolution of directed followership network density across weeks.
Figure S2: Evolution of reciprocity in the directed followership network across weeks.
Figure S3: Average shortest path length in the largest strongly-connected component (panel (a)) and average clustering coefficient (panel (b)) across weeks.
Figure S4: In-degree and out-degree distributions across the year. Each pair of panels illustrates the gradual evolution of the in-degree and out-degree distributions over time.
Appendix G Additional plots on gender scores
Figure S5: Weekly average gender-score with one standard deviation band across all Chirpers. Scores range from 0 to $100$, with lower values reflecting more masculine-coded language and higher values reflecting more feminine-coded language. The mean remains close to the midpoint throughout the year, while the wide standard deviation band indicates substantial heterogeneity in gendered performance every week.
Appendix H Panel regression results for eight-week windows
This appendix reports additional details for the panel regression analysis of social influence described in Section (e) and Appendix D. In particular, we summarise the self and peer coefficients and the instrumental variables diagnostics for the rolling eight-week windows.
Table S4: Panel regression coefficients for eight-week windows: OLS and IV (2SLS) estimates for self-effect ($\varphi$) and peer effect ($\gamma_{\mathrm{inf}}$), both measured in standard deviation units. Statistical significance is indicated as: $^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$, $^{\cdot}p<0.10$; entries labelled (ns) are not significant. The $F$ stat represents the first stage test for the IV model, and the Wu–Hausman $p$ value assesses endogeneity in the OLS results.

### v2

@@ Appendix A Data collection and preprocessing
Table 3: Results of gender score calculation for a subset of a Twitter dataset.
From Table 3, we observe that the model achieved an overall accuracy of 0.84 in both runs. For consistency, we found that 97.4% of profiles (974 out of 1000) were assigned the same gender category in both runs, and 85.4% (854 out of 1000) received the same gender score. These results highlight the model’s robustness in consistently identifying gender across multiple runs.
@@
We use the followership data to construct a sequence of directed networks that summarise the social structure of the Chirpers community over the year. Nodes correspond to Chirpers and a directed edge from $i$ to $j$ indicates that $i$ follows $j$.
@@
Each node in week $t$ is annotated with the gender score $X_{it}$ inferred from its content in that week, when available. If no score is available, meaning that the agent appears in the followership network in week $t$ but has not posted any content in that week, so that $X_{it}$ is missing, we impute its value using the closest available week in time. To do this, we take the most recent previous week when possible and otherwise the nearest subsequent week. These weekly networks, together with the associated gender scores, form the basis for our analysis of gender dynamics, homophily, and the separation between selection and influence in later sections.
Appendix B Scalar assortativity
To quantify homophily based on the continuous gender score, we compute scalar assortativity following Newman (2003), using Equation (21) in [48]. For each week $t$, let $X_{it}$ denote the gender score of agent $i$, and let $E(t)$ denote the set of edges in the followership network. Scalar assortativity is defined as the Pearson correlation between node attributes at the ends of edges,
$r_{t}=\mathrm{corr}(X_{it},X_{jt})_{(i,j)\in E(t)},$
which takes values in the interval [1,1][-1,1]. Positive values indicate assortative mixing (nodes tend to connect to others with similar gender scores), negative values indicate disassortative mixing, and values close to zero indicate no systematic similarity.
Directed assortativity
The Chirper follow network is directed, where a tie $i\rightarrow j$ represents agent $i$ following agent $j$. In the directed case, the edge set is defined as
$E^{dir}(t)=\{(i,j):i\rightarrow j\},$
where $i$ denotes the source node (the follower) and $j$ denotes the target node (the followee). Assortativity, therefore, measures the correlation between the gender score of the source node and that of the target node across all directed edges in week $t$.
Undirected assortativity
For comparison, we also compute assortativity on the network’s undirected projection. In this case, follow ties are treated as undirected edges between agents. The corresponding edge set is defined as
$E^{und}(t)=\{\{i,j\}\},$
where $\{i,j\}$ denotes a pair of nodes connected by at least one follow tie, irrespective of its direction. The assortativity coefficient, therefore, measures similarity in gender scores among connected agents without distinguishing follower and followee roles.
Implementation
All assortativity coefficients are computed using the numeric_assortativity_coefficient function from the NetworkX library. This function implements Newman’s definition of the scalar assortativity coefficient, corresponding to the Pearson correlation of node attributes across edges.
Appendix C Null Networks
@@
For every directed weekly follow graph, we construct $100$ realisations using a directed configuration model that preserves each node’s in-degree and out-degree sequence. Concretely, for week $t$, we draw a directed multigraph from the observed in and out degree sequences, convert it to a simple directed graph by collapsing parallel edges, remove self-loops, relabel nodes back to their original identities, and retain their weekly gender scores $X_{it}$. This yields null networks that match empirical degree heterogeneity while otherwise randomising ties.
@@
We generate an additional $100$ realisations per week using a directed joint degree model, which preserves not only the in and out degree sequences but also the joint distribution of source out-degrees and target in-degrees across edges (the $n_{kk^{\prime}}$ mixing matrix), that is, degree–degree mixing. After removing self-loops and duplicate edges from the empirical weekly graph to estimate $n_{kk^{\prime}}$, we sample null graphs from this joint degree ensemble. We then assign gender scores to null nodes by drawing from the empirical distribution of $X_{it}$ within each $(k_{\mathrm{out}},k_{\mathrm{in}})$ degree class, thereby maintaining any attribute–degree relationship present in the data without imposing additional similarity.
@@
Appendix D STERGM implementation details
@@
Edges are marked as active from their first appearance in the block through the end of the block. In particular, if a directed tie $i\to j$ is first observed in week $t_{0}$ within the block, it is treated as active on $t_{0},T+1)$, where $T$ is the last week in the block. Because follow ties are persistent in our data (we do not observe unfollow events), we do not model dissolution, and the STERGM is fitted in formation-only mode.
@@
Let $V^{\ast}$ denote the global actor roster defined as the union of all identifiers that appear as either followers or followees in any of the 52 weekly edge lists. From these edge lists we recover the weekly gender scores $X_{it}$, where $X_{it}$ denotes the gender score associated with actor $i$ in week $t$. These observations are organised into a long actor–week panel of triples $(t,i,X_{it})$.
Because some actor–week pairs have missing values, we fill short temporal gaps within each actor using last observation carried forward, and next observation carried backward. Actors who never have an observed value in any week are removed, yielding a cleaned roster $\tilde{V}\subseteq V^{\ast}$ and a completed weekly panel of gender scores.
For each week $t$, we define the set of active actors
[…]
Within each block, we restrict attention to actors who appear in at least one edge within the block and intersect this block-specific roster with the cleaned global roster. For each week $t$ in the block, we then extract the completed scores for these actors and standardise them within each week. The resulting standardised scores are attached as a time varying vertex attribute score on the interval $[t,t+1)$ for all actors present in that week.
@@
On each dynamic network, we fit a formation-only STERGM using the tergm package in R [32, 33]. Let $A^{(t)}$ and $A^{(t+1)}$ denote the adjacency matrices at the start and end of an interval, and let $X_{t}$ be the vector of standardised gender scores at time $t$. The formation model specifies, for dyads with $A^{(t)}_{ij}=0$, the conditional probability that a new tie $i\to j$ is formed by $t+1$ as an exponential-family model whose sufficient statistics include an edge formation count, a mutual formation count, and a measure of gender dissimilarity on newly formed ties.
In compact form, the formation component uses the following three effects:
@@
where $s_{\mathrm{edges}}$ controls for baseline formation density, $s_{\mathrm{mutual}}$ captures the tendency to form reciprocal ties, and $s_{\mathrm{abs}}$ accumulates the absolute difference in standardised gender scores across the dyads on which formation is being considered. The corresponding formation parameter vector $\phi=(\phi_{\mathrm{edges}},\phi_{\mathrm{mutual}},\phi_{\mathrm{abs}})^{\top}$ governs the log odds of a new tie forming between weeks $t$ and $t+1$. The coefficient $\phi_{\mathrm{abs}}$ is the primary quantity of interest, with negative values indicating gender-based homophily in the formation of new follow ties and positive values indicating heterophily. We report $\exp(\hat{\phi}_{\mathrm{abs}})$ as the odds ratio for a one-standard-deviation increase in the absolute difference in gendered performance between two actors.
For computational diagnostics, we fit formation-only STERGMs using conditional maximum likelihood estimation with inner models restricted to the maximum pseudo-likelihood (MPLE) step.
Appendix E Panel regression implementation details
@@
[…]
To obtain a complete weekly series of gender scores for each agent, we apply the same forward- and backward-time imputations as in the STERGM construction. Within each agent, we carry the last observed value forward to fill gaps, and then use the next observed value to fill any remaining gaps. Actors who never have an observed gender score in any week are removed from the panel.
Using the followership data, we then compute for each agent $i$ and week $t$ the average gender score of their followees in that week, denoted $\bar{X}_{N(i),t}$. This is obtained by averaging the weekly gender scores of all agents followed by $i$ in week $t$.
For each agent $i$ and week $t\geq 2$ we construct the lagged gender score $X_{i,t-1}$, the lagged mean followee score $\bar{X}_{N(i),t-1}$, and the weekly change
@@
In the final estimation sample we standardise $\Delta X_{it}$, $X_{i,t-1}$ and $\bar{X}_{N(i),t-1}$ to have mean zero and unit variance within each estimation window, so that coefficients can be interpreted in units of standard deviations.
@@
Our main influence model relates weekly changes in gender performance to the agent’s previous week’s gender score and the previous week’s mean gender score of followees:
@@
We first estimate this specification using ordinary least squares (OLS) with agent and week fixed effects, clustering standard errors at the agent level. In this fixed effects OLS model, $\varphi$ measures self-correction, that is, how much an agent’s change in gender score depends on their own previous level after controlling for time-invariant agent heterogeneity and common weekly shocks. The coefficient $\gamma_{\mathrm{inf}}$ measures peer alignment, describing how much an agent’s change depends on the previous week’s mean gender score of their followees. Because all variables are standardised, $\varphi$ and $\gamma_{\mathrm{inf}}$ can be read as changes in $\Delta X_{it}$ in standard deviation units associated with a one standard deviation change in the corresponding predictor.
@@
As an instrument for $\bar{X}_{N(i),t-1}$, we use the mean lagged gender score of followees of followees, that is, agents at distance two in the follow network at time t1t-1. For each week $t$ and agent $i$ we consider all paths of length two of the form
@@
present in the week t1t-1 follow network, where $j$ is a followee of $i$ and $k$ is a followee of $j$. We exclude cases where $k=i$ or where $k$ is already a direct followee of $i$, so the resulting set consists of followees of followees who are not directly followed by $i$. We then take the mean of their lagged gender scores and use this as an instrument $Z_{it}$ for $\bar{X}_{N(i),t-1}$.
In the first stage, we regress the lagged mean followee score on the instrument, including the same agent and week fixed effects as in the outcome equation.
@@
so that $\gamma_{\mathrm{inf}}$ is an instrumental variables estimate of the causal effect of peers on changes in gender performance, under the assumption that followees of followees affect changes in the gender performance of agent $i$ only through their influence on the behaviour of its direct followees.
We use the fixed-effects OLS estimates to describe overall correlations between peer and agent behaviour and the fixed-effects IV estimates to quantify causal peer influence. A one standard deviation increase in the lagged mean gender score of followees is interpreted as producing an estimated $\gamma_{\mathrm{inf}}$ standard deviation change in the agent’s weekly gender score, holding the agent’s previous score, agent effects, and week effects constant.
Windowed estimators
To examine how the strength of influence evolves over time, we estimate the fixed effects OLS and fixed effects IV models on consecutive eight-week windows of the panel, always including an additional look-back week to define the lagged variables required by the model. For each window, we record the estimates of $\varphi$ and $\gamma_{\mathrm{inf}}$ from both OLS and IV specifications, together with standard weak-instrument diagnostics such as the first-stage $F$-statistic and the Wu–Hausman test for endogeneity of the peer term.
Appendix F Structural network measures: detailed definitions
This section provides the formal definitions and computational steps for the structural measures reported in Section 3(a) of the main text. All measures are computed separately for each weekly followership network. Let $A^{(t)}$ denote the directed adjacency matrix for week $t$, and let $n_{t}$ be the number of active agents in that week.
@@ Appendix F Structural network measures: detailed definitions
Social networks are typically sparse, meaning that only a small fraction of all possible ties are present. Tracking density over time reveals how quickly the network becomes more interconnected as agents form follow relationships. Network density measures the proportion of realised ties out of all possible directed ties and is defined as:
@@
Appendix G Additional structural network plots
Figure 6: Evolution of directed followership network density across weeks.
Figure 7: Evolution of reciprocity in the directed followership network across weeks.
Figure 8: Average shortest path length in the largest strongly-connected component (panel (a)) and average clustering coefficient (panel (b)) across weeks.
Figure 9: In-degree and out-degree distributions across the year. Each pair of panels illustrates the gradual evolution of the in-degree and out-degree distributions over time.
Appendix H Additional plots on gender scores
Figure 10: Weekly average gender score with one standard deviation band across all Chirpers. Scores range from 0 to $100$, with lower values reflecting more masculine-coded language and higher values reflecting more feminine-coded language. The mean remains close to the midpoint throughout the year, while the wide standard deviation band indicates substantial heterogeneity in gendered performance every week.
Appendix I Panel regression results for eight-week windows
This appendix reports additional details for the panel regression analysis of social influence described in Section (e) and Appendix E. In particular, we summarise the self and peer coefficients and the instrumental variables diagnostics for the rolling eight-week windows.
Table 4: Panel regression coefficients for eight-week windows: OLS and IV (2SLS) estimates for self-effect ($\varphi$) and peer effect ($\gamma_{\mathrm{inf}}$), both measured in standard deviation units. Statistical significance is indicated as: $^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$, $^{\cdot}p<0.10$; entries labelled (ns) are not significant. The $F$ stat represents the first stage test for the IV model, and the Wu–Hausman $p$ value assesses endogeneity in the OLS results.